OpenStudy (anonymous):

Tracie rolls a die numbered 1 to 6 three times. What is the probability that she will roll a 1, 2, and 3, in that order?

OpenStudy (anonymous):

HELP ME PLEASE

OpenStudy (anonymous):

try dividing six into 100

OpenStudy (anonymous):

I GOT 0.06

OpenStudy (anonymous):

i got 16.67

OpenStudy (anonymous):

6 DIVIDE 100=0.06

OpenStudy (anonymous):

100 divided by 6

OpenStudy (anonymous):

16.66

OpenStudy (anonymous):

16.67 because the six behind the decimal is a run on intil it reaches seven

OpenStudy (anonymous):

O OK

OpenStudy (anonymous):

so its 1/36

OpenStudy (klimenkov):

\(\frac16\cdot\frac16\cdot\frac16\)

OpenStudy (anonymous):

so its 1/36

OpenStudy (klimenkov):

No.

OpenStudy (anonymous):

1/216?

OpenStudy (anonymous):

1/24

OpenStudy (klimenkov):

\(\frac{1}{216}\)

OpenStudy (anonymous):

i new it!

OpenStudy (anonymous):

i just learned something new today, i can be wrong sometimes

OpenStudy (anonymous):

ILOVEYOUTOBY sorry